OpenStudy (anonymous):

Solve using the square root method. 9x2 - 25 = 0

OpenStudy (anonymous):

\[\LARGE 9=3^2\quad \quad , \quad \quad 25=5^2\] \[\LARGE (3x)^2-5^2=0\] now according to... \[\LARGE a^2-b^2=(a-b)(a+b)\] you have: \[\LARGE (3x-5)(3x+5)=0\] from here you have: 3x-5=0 or 3x+5=0 solve them.. you'll have the answer ;)

OpenStudy (anonymous):

I get 1.6 and -1.6. Is that right?

OpenStudy (anonymous1):

Yes, that is correct. 1.6666... and -1.6666...